OpenStudy (anonymous):

1. x – 3 > –9 2. –2x + 1 ≤ –11 3. 10 < –3x + 1 4. 2(x + 5) > 8x – 8 5. –2(x – 3) ≥ 5 – (x + 3)

OpenStudy (anonymous):

cmon man don't put all your questions, did you at least try solving any?

OpenStudy (anonymous):

I need help I don't expect all the answers I just need help!

OpenStudy (anonymous):

Just do it like you would normally don't let the signs scare you, but remember when you divide by a negative sign you switch the sign (greater than, less than, etc.)

OpenStudy (anonymous):

1.\[x-3>-9\]\[x>-6\] 2.\[-2x+1 \le -11\]\[-2x \le -12\]\[x \ge 6\] 3. \[10<-3x+1\]\[9 < -3x\]\[x < -3\] 4.\[2(x+5) >8x-8\]\[2x+10>8x-8\]\[-6x>-18\]\[x<3\] 5.\[–2(x – 3) ≥ 5 – (x + 3)\]\[-2x+6 \ge 5-x-3\]\[-x \ge -4\]\[x \le 4\]